Number 9?
Use the following information/data to answer questions 8 and 9: ou have invented a new biodegradable chemical spray that will protect developing almond nuts from seed eating weevil insects. In a clinical trial, the almonds on one side of an almond tree were sprayed with your chemical while th Treatment Type Chemical Spray Controls e other side almonds were left untreated to serve as the controls. The results are as follows: # of seeds with weevils % seeds weevil infected # of almond seeds 245 213 182 74.2857143% 191 89.6713615% 8. Calculate the 95% confidence interval for the percentage (proportion) of sprayed seeds that were weevil infected. Remember, z 1.960: a.) Standard Method b.) Wilson Method c.) Safe/Conservative Method 9. Calculate the 95% confidence interval for the difference in percentage of seeds that were infected for these two treatments using the standard method ANDİNTERPRETTHISCONFIDENCEINTERVAL. Remember, z = 1.960.
Answer #1

pi-182/245 P2-191/213 0.74286 0.89671 75.286% 89.671%

The 95% per cent confidence interval for p1-p2 is given by:
(\hat{p}_1-\hat{p}_2)\pm z_{\alpha/2} \sqrt{\dfrac{\hat{p}_1(1-\hat{p}_1)}{n_1}+\dfrac{\hat{p}_2(1-\hat{p}_2)}{n_2}}

pı and fz are the proportions of two groups.

z_{\alpha/2}is the critical value
             

\sqrt{\dfrac{\hat{p}_1(1-\hat{p}_1)}{n_1}+\dfrac{\hat{p}_2(1-\hat{p}_2)}{n_2}} is the SE of the difference

So confidence interval is

(0.89671-0.74286)\pm 1.96* \sqrt{\dfrac{0.74286(1-0.74286)}{245}+\dfrac{0.89671(1-0.89671)}{213}}(0.085544,0.22215)

or (8.5544\%,22.215\%)

So 95% confidence interval for the difference in the percentage of seeds that were infected for these two treatments is between 8.5to 22.2 %